# A System Level Approach to LQR Control of the Diffusion Equation

### v1

@@ III Problem Set Up
$\rho(A)$ will contain a set of the form $\{s\ \in\mathbb{C};~{\rm Re}(s)\geq\omega\}$, where $\omega$ defines a growth bound on the semigroup $\{T_{A}(t)\}$. When this semigroup is exponentially stable, one can take $\omega=0$ [14, Ch. 2]. Then the operator-valued transfer function representation of $\Sigma(A,B,C,D)$ is given by
@@
We say this transfer function is strictly proper if $\lim_{s\rightarrow\infty}\left(C(sI-A)^{-1}B+D\right)=0,$ i.e. if $D=0$.
When $\Sigma(A,B,C,D)$ is spatially invariant, each of $A,B,C,D$ can be transformed to a multiplication operator on $\ell^{2}(\mathbb{Z})$ spaces of appropriate dimensions, e.g., $M_{\hat{A}}=\mathcal{F}^{-1}A\mathcal{F}$. In this case, (17) can be represented as
$\tfrac{d}{dt}\hat{{\boldsymbol{\psi}}}(t)$ $=(M_{\hat{A}}\hat{{\boldsymbol{\psi}}})(t)+(M_{\hat{B}}\hat{\bf u})(t)$ (13)
$\hat{\bf z}(t)$ $=(M_{\hat{C}}\hat{{\boldsymbol{\psi}}})(t)+(M_{\hat{D}}\hat{\bf u})(t),$
and its transfer function can also be represented as a multiplication operator [14, Ch.7]:
$\mathcal{F}^{-1}\left(C(sI-A)^{-1}B+D\right)\mathcal{F}=M_{\hat{G}(s)}$ (14)
$\hat{G}_{\kappa}(s)=\hat{C}_{\kappa}(sI-\hat{A}_{\kappa})^{-1}\hat{B}_{\kappa}+\hat{D}_{\kappa}.$
III-A Controller Design for the Diffusion Equation
We focus here on a specific plant - the diffusion equation - noting that our approach could be generalized to other linear distributed parameter systems. The diffusion equation over $\mathbb{T}$ with fully distributed actuation $u$ and disturbance $w$ is described by the partial differential equation (PDE)
where $\alpha>0$ is the constant of diffusivity. The spatio-temporal signal
${z}(t,\theta)=\left[\begin{array}[]{c}\gamma{\psi}(t,\theta)\\ {u}(t,\theta)\end{array}\right],$ (16)
specifies a performance output of interest that provides a uniform penalty across the spatial variable, with $\gamma>0$ determining a tradeoff between cost of state and cost of control. The dynamics (15)-(16) can be written as
$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A{\boldsymbol{\psi}})(t)+{\bf u}(t)+{\bf w}(t)$ (17)
${\bf z}(t)$ $=(C{\boldsymbol{\psi}})(t)+(D{\bf u})(t),$
where $A\in\mathcal{L}(L_{1}^{2}(\mathbb{T}))$ is defined by
$(A\psi)(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)$ (18)
and generates a $C_{0}$- semigroup $\{T_{A}(t)\}$ on
with derivatives, $f^{\prime},f^{\prime\prime}$, in (19) interpreted weakly, and $C,D\in\mathcal{L}(L_{1}^{2}(\mathbb{T}),L_{2}^{2}(\mathbb{T}))$ are the bounded operators
$(C\psi)(t,\theta)=\left[\begin{array}[]{c}\gamma\psi(t,\theta)\\ 0\end{array}\right],~(Du)(t,\theta)=\left[\begin{array}[]{c}0\\ u(t,\theta)\end{array}\right].$ (20)
We denote the operator-valued transfer function for (17) by
$P(s)=C(sI-A)^{-1}+D,$ (21)
which is well-defined on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\subset\rho(A)$ since 0 defines a growth bound on the semigroup $\{T_{A}(t)\}.$ $\Sigma(A,I,C,D)$ is a spatially-invariant system that is equivalently characterized by its transformation under $\mathcal{F}$ as
$\tfrac{d}{dt}\hat{\psi}(t,\kappa)=\underbrace{-\alpha\kappa^{2}}_{\hat{A}_{\kappa}}\hat{\psi}(t,\kappa)+\hat{u}(t,\kappa)+\hat{w}(t,\kappa),$ (22)
$\hat{z}(t,\kappa)=\underbrace{\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right]}_{\hat{C}_{\kappa}}\hat{\psi}(t,\kappa)+\underbrace{\left[\begin{array}[]{c}0\\ 1\end{array}\right]}_{\hat{D}_{\kappa}}\hat{u}(t,\kappa),~~t\in[0,\infty),~\kappa\in\mathbb{Z}$
or in transfer function form as $\mathcal{F}^{-1}P(s)\mathcal{F}=M_{\hat{P}(s)},$
$\hat{P}_{\kappa}(s)=\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right](s+\alpha\kappa^{2})^{-1}+\left[\begin{array}[]{c}0\\ 1\end{array}\right].$ (23)
We consider a state feedback control policy defined by a transfer function
@@ III-A Controller Design for the Diffusion Equation
defined on some set $\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$. The feedback interconnection of (17) and (24) is a well-defined transfer function on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\cap\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$ [14, Ex. 7.3], which we denote by $\mathbb{F}(P;K)(s).$
Definition 2
We say that the control policy (24) internally stabilizes the system (17) if $\mathbb{F}(P;K)(s)$ is the operator-valued transfer function of a system that generates an exponentially stable $C_{0}$-semigroup $\{T_{A_{\rm c.l.}}(t)\}$ and $\|\mathbb{F}(P;K)\|_{H_{2}}<\infty$.
We now state formally the $H_{2}$ design problem of interest:
$\inf_{K}~~\|\mathbb{F}(P;K)\|_{H_{2}}$ (25)
$~{\rm s.t.}~~K~\text{spatially-invariant}$
$~~~~~~K~\text{internally stabilizing}.$
As in the finite-dimensional setting, this $H_{2}$ problem will recover the solution to the Linear Quadratic Regulator (LQR) problem for the diffusion equation [15, 14]:
$\inf_{u}~\int_{0}^{\infty}\left<{\boldsymbol{\psi}}(t),\gamma^{2}{\boldsymbol{\psi}}(t)\right>_{L^{2}(\mathbb{T})}+\left<{\bf u}(t),{\bf u}(t)\right>_{L^{2}(\mathbb{T})}dt$ (26)
[…]
IV Closed-Loop Formulation of $H_{2}$ Problem
In this section, we demonstrate that the controller design problem (LABEL:eq:H2_formulation) can be formulated as an optimization problem over certain closed-loop mappings and admits analytical solutions in the spatial frequency domain. We will illustrate that this formulation allows for spatio-temporal constraints to be imposed in a convex manner in Section VII.
We isolate the closed-loop transfer functions from $w$ to $\psi$ and $u$ in the interconnection of (17) and (24) to obtain:
@@ IV Closed-Loop Formulation of $H_{2}$ Problem
The operator-valued transfer functions $\Phi^{\psi}(s),\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ are analogous to the “system-level maps” introduced for finite dimensional discrete-time systems in [5]. The closed-loop transfer function from $w$ to $z$ can be written in terms of $\Phi^{\psi}(s),\Phi^{u}(s)$ as
@@
See Appendix -C. ∎
Given the analogy to the definition for finite dimensional systems, we say that an operator-valued transfer function is an element of $H_{\infty}$ if it satisfies the condition (LABEL:eq:Hinf_condition).
@@
Let $\Phi^{\psi}(s)$ and $\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ be defined by (27) for some spatially-invariant and internally stabilizing controller $K(s).$ Then, for all $\eta,\vartheta\in L^{2}(\mathbb{T})$ and $s\in\{s\in\mathbb{C};{\rm Re}(s)\geq 0\}$,
@@
See Appendix -D. ∎
@@
Let ${\Phi}^{\psi},{\Phi}^{u}\in H_{\infty}\subset\mathcal{L}(L^{2}(\mathbb{T}))$ be spatially-invariant, strictly proper, and satisfy (32). Select a spatially-invariant transfer function [^1] Existence of such a $G$ will be verified in Section V-A. $G$ for which $G^{-1}(s)\in H_{\infty}$, $\lim_{s\to\infty}s\hat{G}_{\kappa}^{-1}(s)=1~\forall\kappa\in\mathbb{Z}$, and $G^{-1}(s)(sI-A)\in H_{\infty}$. Then,
@@
defines an internally stabilizing controller for (17). This implementation (33) is well-posed and the internal signal $\bf v$} remains bounded for all bounded signals $\bf n$} and $\bf w$} as depicted in Figure 1. The input-output map (transfer function) corresponding to (33) is
${\bf U}=\mathcal{F}^{-1}M_{\hat{K}(s)}\mathcal{F}{{\boldsymbol{\Psi}}},~~\hat{K}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}.$ (34)
@@
Lemmas 1-3 demonstrate that the set of all internally stabilizing and spatially-invariant controllers for (17) can be characterized by an affine subspace of spatially-invariant transfer functions $\Phi^{\psi},\Phi^{u}.$ With this, we are able to state our main result.
Theorem 1
The following is equivalent to the $H_{2}$ problem of interest (LABEL:eq:H2_formulation) for the diffusion equation:
$\inf_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{H_{2}}$ (35)
[…]
V Analytical Solution to Closed-Loop Formulation of $H_{2}$ Problem
Spatial invariance properties allow us to solve (LABEL:eq:CL_formulation) in the spatial frequency domain. By the Placherel theorem, formulation (LABEL:eq:CL_formulation) of the $H_{2}$ design problem for the diffusion equation is equivalent to the optimization problem:
$\inf_{{\hat{\Phi}^{\psi}(s)},{\hat{\Phi}^{u}(s)}}~\left\|M_{\hat{C}}M_{\hat{\Phi}^{\psi}(s)}+M_{\hat{D}}M_{\hat{\Phi}^{u}(s)}\right\|_{H_{2}}$ (36)
[…]
Note that the cost and constraints of (LABEL:eq:CLOpt_Frequency) “decouple” in the spatial frequency variable $\kappa.$ Thus, the solution to (LABEL:eq:CLOpt_Frequency) can be obtained by solving the family of $H_{2}$ problems, parameterized by $\kappa\in\mathbb{Z}:$
$\inf_{\hat{\Phi}^{\psi}_{\kappa},\hat{\Phi}^{u}_{\kappa}}\left\|\left[\begin{array}[]{cc}\gamma&0\\ 0&1\end{array}\right]\left[\begin{array}[]{c}\hat{\Phi}^{\psi}_{\kappa}\\ \hat{\Phi}^{u}_{\kappa}\end{array}\right]\right\|_{H_{2}}$ (37e)
$~~{\rm s.t.}~~~\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}\in RH_{2}$ (37f)
$~~~~~~~~\left((s+1)-(-\alpha\kappa^{2}+1)\right)\hat{\Phi}^{\psi}_{\kappa}(s)-\hat{\Phi}_{\kappa}^{u}(s)=1,$ (37g)
provided that the family of solutions $\{\hat{\Phi}^{\psi}_{\kappa}\},~\{\hat{\Phi}^{u}_{\kappa}\}$ satisfy the condition (LABEL:eq:Hinf_frequency) which ensures $\Phi^{\psi},\Phi^{u}\in H_{\infty}$.
For each fixed $\kappa\in\mathbb{Z},$ (37) is a standard $H_{2}$ design problem for a finite-dimensional system with an implicit affine subspace constraint (37g). To obtain an analytical solution to this problem, we first replace the implicit constraint (37g) with an analogous explicit constraint.
Lemma 4
Fix $\kappa\in\mathbb{Z}.$ The single-input single-output (SISO) transfer functions $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ satisfy $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}\in RH_{2}$ and (37g) if and only if they are of the form
$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\frac{1}{s+1}\left(1+\hat{\rho}_{\kappa}(s)\right),$ (38)
$\hat{\Phi}^{u}_{\kappa}(s)$ $=\frac{\alpha\kappa^{2}-1}{s+1}+\frac{s+\alpha\kappa^{2}}{s+1}\hat{\rho}_{\kappa}(s)$
for some SISO transfer function $\hat{\rho}_{\kappa}(s)\in RH_{2}.$
Proof:
See Appendix -F. ∎
We leverage the parameterization (38) and properties of the $H_{2}$ norm to write (37) as an unconstrained $H_{2}$ design problem over a single SISO transfer function decision variable $\hat{\rho}_{\kappa}.$
Proposition 1
The optimal solution $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ to the problem (37) can be recovered from the solution $\hat{\rho}_{\kappa}$ to the optimization problem
$\inf_{\hat{\rho}_{\kappa}\in RH_{2}}~\left\|\left[\begin{array}[]{c}\tfrac{\gamma}{s+1}\\ \tfrac{\alpha\kappa^{2}-1}{s+1}\end{array}\right]+\left[\begin{array}[]{c}\tfrac{\gamma}{s+1}\\ \tfrac{s+\alpha\kappa^{2}}{s+1}\end{array}\right]\hat{\rho}_{\kappa}(s)\right\|_{H_{2}}$ (39)
through the relation (38). The optimal solution to (39) is
$\hat{\rho}_{\kappa}^{\rm opt.}(s)=\frac{1-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}},$ (40)
and thus the optimal solution to (37) is given by
$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\frac{1}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}},$ (41)
Proof:
See Appendix -A. ∎
V-A Implementation of Controller via Closed-Loop Solution
@@
to satisfy all conditions of Lemma 3. In this case, (33) may be specified in spatial frequency $\kappa$ as
@@
each of which is verified to be the symbol of a multiplication operator in $H_{\infty}$ for our selected $\hat{G}$. We numerically compute the spatio-temporal impulse responses (Green’s functions) of the blocks in Fig. 1 and plot their dynamic and static components in Fig. 2. We note that the kernels of the dynamic components are “more localized” for smaller time values and appear to spread out spatially as time increases.

### v2

@@ IV-A Transfer Functions
$\rho(A)$ will contain a set of the form $\{s\ \in\mathbb{C};~{\rm Re}(s)\geq\omega\}$, where $\omega$ defines a growth bound on the semigroup $\{T_{A}(t)\}$. When this semigroup is exponentially stable, one can take $\omega=0$ [7, Ch. 2]. Then the operator-valued transfer function representation of a system (5) $\Sigma(A,B,C,D)$ is given by
@@
If $A$ generates an exponentially stable $C_{0}$-semigroup, then we say that the transfer function representation of $\Sigma(A,B,C,D)$ is exponentially stable. If $D\equiv 0$, the transfer function representation of $\Sigma(A,B,C,D)$ is strictly proper.
Now let $\Sigma(A,B,C,D)$ describe the diffusion equation (13). We denote the operator-valued transfer function from inputs $\begin{bmatrix}\boldsymbol{w}&\boldsymbol{u}\end{bmatrix}^{\top}$ to $\begin{bmatrix}\boldsymbol{z}&\boldsymbol{\psi}\end{bmatrix}^{\top}$ for (13) by $P(s)$, which is well-defined on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\subset\rho(A)$ as 0 defines a growth bound on the semigroup $\{T_{A}(t)\}.$ Since (13) is a spatially invariant system, $P(s)$ can be expressed as a multiplication operator, $M_{\hat{P}(s)},$ over spatial frequency [7, Ch.7], and (LABEL:eq:open_loop_kappa) can be represented by the transfer function
$\hat{P}_{\kappa}(s)=\begin{bmatrix}\hat{C}_{\kappa}(s+\alpha\kappa)^{-1}&\hat{C}_{k}(s+\alpha\kappa)^{-1}+\hat{D}_{\kappa}\\ (s+\alpha\kappa)^{-1}&(s+\alpha\kappa)^{-1}\end{bmatrix}$, (20)
which maps $\begin{bmatrix}\hat{w}_{\kappa}(s)&\hat{u}_{\kappa}(s)\end{bmatrix}^{\top}$ to $\begin{bmatrix}\hat{z}_{\kappa}(s)&\hat{\psi}_{\kappa}(s)\end{bmatrix}^{\top}$.
We consider a (possibly dynamic) state feedback control policy defined by a spatially invariant transfer function
@@
defined on some set $\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$. The feedback interconnection of $P(s)$ and (21), denoted $\mathbb{F}(P;K)(s)$, is a well-defined transfer function on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\cap\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$ [7, Ex. 7.3] that maps $\boldsymbol{w}(s)$ to $\boldsymbol{z}(s)$.
IV-B Closed-Loop Mappings
We isolate the closed-loop transfer functions from $\boldsymbol{w}$} to $\boldsymbol{\psi}$} and $\boldsymbol{u}$} in $\mathbb{F}(P;K)(s)$ to obtain:
@@ IV-B Closed-Loop Mappings
The operator-valued transfer functions $\Phi^{\psi}(s),\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ for each $s\in\mathbb{C}$ at which they are defined and are analogous to the “system-level maps” introduced for finite dimensional systems in [22]. The closed-loop transfer function from $\boldsymbol{w}$} to $\boldsymbol{z}$} can be written in terms of $\Phi^{\psi}(s),\Phi^{u}(s)$ as
@@
See Appendix -B. ∎
Given the analogy to the definition for finite dimensional systems, we say that an operator-valued transfer function is an element of $\mathcal{H}_{\infty}$ if it satisfies the condition (LABEL:eq:Hinf_condition).
@@
Let $\Phi^{\psi}(s)$, $\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ be defined by (22) for some spatially-invariant and internally stabilizing controller $K(s).$ Then, for all $\eta,\vartheta\in L^{2}(\mathbb{T})$ and $s\in\{s\in\mathbb{C};{\rm Re}(s)\geq 0\}$,
@@
See Appendix -C. ∎
We can now state our main result:
Theorem 1
The solution to the $\mathcal{H}_{2}$ design problem for the diffusion equation (17) can be recovered from the solution to:
$\operatorname*{arg\,min}_{\Phi^{\psi},\Phi^{u}\in\mathcal{H}_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{\mathcal{H}_{2}}$ (28)
[…]
Lemmas 1-2 show that $K$ spatially invariant and internally stabilizing implies that $\Phi^{\psi},\Phi^{u}\in\mathcal{H}_{\infty}$ and are spatially invariant, strictly proper, and satisfy the affine constraint (LABEL:eq:affine_innerProduct). To complete the proof of Theorem 1 it remains to show that any such $\Phi^{\psi},\Phi^{u}$ can be implemented as an internally stabilizing controller. We will prove this in Lemma 3 and discuss controller implementation in Section V.
IV-C Analytical Solution to Closed-Loop Formulation of $\mathcal{H}_{2}$ Problem
Spatial invariance allow us to solve (LABEL:eq:CL_formulation) in the spatial frequency domain. By the Plancherel theorem, formulation (LABEL:eq:CL_formulation) of the $\mathcal{H}_{2}$ design problem for the diffusion equation is equivalent to the optimization problem:
$\operatorname*{arg\,min}_{{\hat{\Phi}^{\psi}(s)},{\hat{\Phi}^{u}(s)}}~\left\|M_{\hat{C}}M_{\hat{\Phi}^{\psi}(s)}+M_{\hat{D}}M_{\hat{\Phi}^{u}(s)}\right\|_{\mathcal{H}_{2}}$ (29)
[…]
Note that the cost and constraints of (LABEL:eq:CLOpt_Frequency) “decouple” in the spatial frequency variable, so the solution to (LABEL:eq:CLOpt_Frequency) can be obtained by solving a parameterized family of $\mathcal{H}_{2}$ problems for finite-dimensional systems parameterized by $\kappa\in\mathbb{Z}$. Let ${\mathcal{RH}}_{2}$ denote the set of rational transfer functions with strictly proper entries and no poles in the closed right half plane [8, Ch. 3]. Then the symbol of the minimizer of (LABEL:eq:CLOpt_Frequency) at frequency $\kappa$ is the solution to the family of problems:
$\operatorname*{arg\,min}_{\hat{\Phi}^{\psi}_{\kappa},\hat{\Phi}^{u}_{\kappa}\in\mathcal{RH}_{2}}\left\|\left[\begin{array}[]{cc}\gamma&0\\ 0&1\end{array}\right]\left[\begin{array}[]{c}\hat{\Phi}^{\psi}_{\kappa}\\ \hat{\Phi}^{u}_{\kappa}\end{array}\right]\right\|_{\mathcal{H}_{2}}$ (30e)
$~~{\rm s.t.}\left((s+1)-(-\alpha\kappa^{2}+1)\right)\hat{\Phi}^{\psi}_{\kappa}(s)-\hat{\Phi}_{\kappa}^{u}(s)=1,$ (30f)
provided that the family of solutions $\{\hat{\Phi}^{\psi}_{\kappa}\},~\{\hat{\Phi}^{u}_{\kappa}\}$ satisfy the condition (LABEL:eq:Hinf_frequency) which ensures $\Phi^{\psi},\Phi^{u}\in\mathcal{H}_{\infty}$.
For each fixed $\kappa\in\mathbb{Z},$ (30) is a standard $\mathcal{H}_{2}$ design problem for a finite-dimensional system with an implicit affine subspace constraint (30f), which we can solve analytically.
Proposition 1
The optimal solution $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ to the problem (30), with $\gamma$ defined in (12), is given by
$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\left(s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}\right)^{-1},$ (31)
Proof:
See Appendix -A. ∎
V internally stable implementation of Controller via Closed-Loop Solution
We now complete the proof of Theorem 1 by showing that $\Phi^{\psi},\Phi^{u}$ satisfying the constraints of (LABEL:eq:CL_formulation) can be used to construct an internally stable implementation of the controller, which thus preserves the structural properties of these maps. In order for an implementation of the controller to be well posed and preserve the structure of $\Phi^{\psi}$ and $\Phi^{u}$, we introduce an additional transfer function $G$ that ensures internal states of the controller remain bounded.
@@ IV-C Analytical Solution to Closed-Loop Formulation of $\mathcal{H}_{2}$ Problem
Let ${\Phi}^{\psi},{\Phi}^{u}\in\mathcal{H}_{\infty}\subset\mathcal{L}(L^{2}(\mathbb{T}))$ be spatially-invariant, strictly proper, and satisfy (27). Select a spatially-invariant transfer function $G$ for which $G^{-1}(s)\in\mathcal{H}_{\infty}$, $\lim_{s\to\infty}s\hat{G}_{\kappa}^{-1}(s)=1~\forall\kappa\in\mathbb{Z}$, and $G^{-1}(s)(sI-A)\in\mathcal{H}_{\infty}$. Then,
@@
defines an internally stable implementation of the controller in the sense that the closed loop transfer matrices from the exogenous inputs $(\hat{n}_{\kappa}(s),\hat{w}_{\kappa}(s))$ to the internal signals $(\hat{\psi}_{\kappa}(s),\hat{u}_{\kappa}(s),\hat{v}_{\kappa}(s))$ belong to $\mathcal{H}_{\infty}$. This implementation is depicted in Figure 1. The input-output map (transfer function) corresponding to (32) is
${\bf U}=\mathcal{F}^{-1}M_{\hat{K}(s)}\mathcal{F}{{\boldsymbol{\Psi}}},~~\hat{K}_{\kappa}(s)=\tfrac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}.$ (33)
@@
to satisfy all conditions of Lemma 3. Note that this is one suitable choice of $G$, but it is not obvious how to “best” select $G$ in general. Then, (32) is specified at frequency $\kappa$ as
@@
each of which is verified to be the symbol of a multiplication operator in $\mathcal{H}_{\infty}$ for our selected $\hat{G}$. We numerically compute the spatio-temporal impulse responses (Green’s functions) of the blocks in Fig. 1 and plot their dynamic and static components over the spatial domain for various selections of time in Fig. 2. We note that the kernels of the dynamic components are “more localized” for smaller time values and appear to spread out spatially as time increases.